OpenStudy (kevin):

value of cos x - sqrt(3) sin x > 0, if ....

OpenStudy (kevin):

@3mar @zepdrix

OpenStudy (kevin):

@agent0smith

OpenStudy (kevin):

The answer is \[\frac{ \pi }{ 12 } < x < \frac{ \pi }{ 6 }\] Teach me how to get that

OpenStudy (kevin):

@Kainui

OpenStudy (3mar):

Hello!

OpenStudy (3mar):

How are you?

OpenStudy (kevin):

I'm fine :) u?

OpenStudy (3mar):

I am fine. Thank you!

OpenStudy (kevin):

glad to hear that! can u help me solve this problem? I need to solve it quickly

OpenStudy (kevin):

@FaiqRaees

OpenStudy (3mar):

Let me try

OpenStudy (kevin):

ok thx :)

OpenStudy (faiqraees):

Is this the problem \[ \cos x - \sqrt[3]{\sin x}>0 \]

OpenStudy (kevin):

\[\cos x - \sqrt{3} \sin x > 0\]

OpenStudy (faiqraees):

oh okay start by taking the sin x term on right side

OpenStudy (3mar):

Yes it looks like what Kevin has just typed

OpenStudy (kevin):

well, like what?

OpenStudy (3mar):

tp be like that \[\cos(x)>\sqrt{3}\sin(x)\]

OpenStudy (3mar):

*to

OpenStudy (kevin):

then?

OpenStudy (faiqraees):

Divide sqrt3 cos x on both sides

OpenStudy (3mar):

then divide the two sides by cos(x)

OpenStudy (3mar):

\[\frac{ \cos(x) }{ \cos(x) }>\sqrt{3}\frac{ \sin(x) }{ \cos(x) }\]

OpenStudy (kevin):

1 > sqrt(3) tan x

OpenStudy (faiqraees):

divide sqrt 3 on both sides

OpenStudy (3mar):

to be: \[1>\sqrt{3}\tan(x)\] finally you got the form: \[\tan(x)<\frac{ 1 }{ \sqrt{3} }\]

OpenStudy (kevin):

ok....

OpenStudy (kevin):

Idk but, my book said tan x = -sqrt 3

OpenStudy (3mar):

What is the domain we can get the answer in?

OpenStudy (kevin):

there is no domain

OpenStudy (3mar):

Can you share the answer procedure?

OpenStudy (3mar):

of your book?

OpenStudy (kevin):

it's too long

OpenStudy (kevin):

wait

OpenStudy (3mar):

Take your time.

OpenStudy (3mar):

To be aware of what the result be, this is a graph of it. sorry it will be that: https://www.desmos.com/calculator/9yohdvzrq3

OpenStudy (kevin):

pls wait

OpenStudy (faiqraees):

Your book answer is wrong. Take for example the angle -23. The tan of -23 is -0.424 which is smaller than 1/sqrt 3. And the cos (-23) - sqrt3 sin(-23) = 1.59

OpenStudy (faiqraees):

23 degrees ~ -2/15 radians

OpenStudy (3mar):

Let's check firstly, FaiqRaees

OpenStudy (kevin):

http://prnt.sc/cuw3pn

OpenStudy (faiqraees):

How did the book interchange between inequality sign and equal to sign in the fifth step?

OpenStudy (kevin):

it's not done yet. here is the second page : http://prnt.sc/cuw5ej

OpenStudy (kevin):

I will translate it

OpenStudy (faiqraees):

No Okay I got it. I made a slight easier approach wanna hear?

OpenStudy (kevin):

btw I always lost connection, I must refresh everytime I post something here. Ok I glad to hear that

OpenStudy (faiqraees):

Okay we know that cos pi/2 and cos 3pi/2 =0 Which means we can write cos (x-2pi/3) = cos pi/2 And cos (x-2pi/3) = cos 3pi/2 Right?

OpenStudy (kevin):

where did you get cos pi/2 and cos 3pi/2 ?

OpenStudy (faiqraees):

take the arc cos of 0. You will get pi/2 and 3pi/2

OpenStudy (kevin):

Sorry, but where did you get arc cos of 0?

OpenStudy (faiqraees):

Can you draw the graph of cos x

OpenStudy (kevin):

I mean on the question, it's not said that cos x = 0

OpenStudy (faiqraees):

Oh that. We are taking the extreme values possible. By that we can work out the boundaries

OpenStudy (faiqraees):

Its like if I say what are the values of x for which f(x)>0. By working out the value of x at which f(x)= 0, you can easily tell for which values of x is f(x)>0 (Considering f(x) to be increasing function)

OpenStudy (kevin):

Hmm.... ok, I'll study it further xD continue your answer explanation

OpenStudy (kevin):

How about this cos (x-2pi/3) = cos pi/2 And cos (x-2pi/3) = cos 3pi/2 Where did you get that?

OpenStudy (faiqraees):

Can you understand till the fifth step in your book

OpenStudy (kevin):

no xD

OpenStudy (kevin):

how about u?

OpenStudy (faiqraees):

Well they first convert the expression into the form kcos (x-a) Clear?

OpenStudy (kevin):

no

OpenStudy (kevin):

I don't see the relation

OpenStudy (kevin):

o_O

OpenStudy (faiqraees):

wait 2 minutes.

OpenStudy (kevin):

ok

OpenStudy (kevin):

@3mar any idea?

OpenStudy (faiqraees):

Btw you dont understand how kcos(x-a) was arrived?

OpenStudy (kevin):

it's from cos x = cos a

OpenStudy (kevin):

am I right? ._.

OpenStudy (faiqraees):

No no. You haven't studied the angle addition formula?

OpenStudy (kevin):

Owhh... I got it, it's from a cos x + b sin x I never teached by my teacher about different trig sum

OpenStudy (kevin):

if u know what I mean

OpenStudy (faiqraees):

yes its from that

OpenStudy (kevin):

then how did second line on my book come from?

OpenStudy (kevin):

I mean third line- fourth

OpenStudy (faiqraees):

Yeah the answer in your book is wrong https://www.wolframalpha.com/input/?i=cos+x+-sqrt3+sin+x+%3E0

OpenStudy (faiqraees):

For the answer to be pi/12 and pi/6, the value of n in the real solution has to be 11/24 and since n belongs to the set of integer it cannot have the value 11/24

OpenStudy (kevin):

But, your answer is not provided on the options, is it?

OpenStudy (faiqraees):

No the answer I mentioned pi/12 and pi/6 were given by your book

OpenStudy (kevin):

Owhh... ok, anyway let's using my book way cos I learned from it xD I've know line 1-4 now. How about fifth line?

OpenStudy (kevin):

This is a question from a test that given by my government to selected the student who want to go a college. That's why it's not teached intensively in my school.

OpenStudy (kevin):

I'm understand now, thx guys!

OpenStudy (agent0smith):

value of cos x - sqrt(3) sin x > 0, if .... if what??? Where's the rest of the question??

OpenStudy (agent0smith):

@FaiqRaees @3mar you should not divide both sides by cos x, since you do not know if it's positive. You can't do it w/o affecting the truthiness of the inequality.

OpenStudy (kevin):

That's the question @agent0smith I've figured out the answer, thx for u reply :)